# Quantum Blackwell’s Ordering and Differential Privacy

### v1

@@ II Notations, Definitions and Facts
Proof.
We first prove that (1) implies (2). From (25), we have that for any $0\preceq\Lambda\preceq\mathbb{I}$,
$\mathrm{Tr}[\Lambda\rho]-e^{\varepsilon}\mathrm{Tr}[\Lambda\sigma]\leq\delta,$
which can be rewritten as
$\mathrm{Tr}[\Lambda(\rho-e^{\varepsilon}\sigma)]\leq\delta.$
Taking the maximum over all $0\preceq\Lambda\preceq\mathbb{I}$ on the left-hand side, we get
$\max_{0\preceq\Lambda\preceq\mathbb{I}}\mathrm{Tr}[\Lambda(\rho-e^{\varepsilon}\sigma)]\leq\delta.$
By the definition of the quantum hockey stick divergence (Definition 13), the left-hand side is exactly $E_{e^{\varepsilon}}(\rho\|\sigma)$. Thus, $E_{e^{\varepsilon}}(\rho\|\sigma)\leq\delta$.
Next, we prove that (2) implies (1). Assume $E_{e^{\varepsilon}}(\rho\|\sigma)\leq\delta$. By definition, this means
$\max_{0\preceq\Lambda^{\prime}\preceq\mathbb{I}}\mathrm{Tr}[\Lambda^{\prime}(\rho-e^{\varepsilon}\sigma)]\leq\delta.$
This implies that for any specific choice of $0\preceq\Lambda\preceq\mathbb{I}$, we must have
$\mathrm{Tr}[\Lambda(\rho-e^{\varepsilon}\sigma)]\leq\max_{0\preceq\Lambda^{\prime}\preceq\mathbb{I}}\mathrm{Tr}[\Lambda^{\prime}(\rho-e^{\varepsilon}\sigma)]\leq\delta.$
Rearranging the terms, we get
$\mathrm{Tr}[\Lambda\rho]\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\sigma]+\delta,$
which is the condition in (25). This completes the proof.
Fact 19 ([31, Lemma $6.9$]).
If $E_{e^{\varepsilon}}(\rho\|\sigma)\leq\delta$, for a pair $(\rho,\sigma)\in\mathcal{D}(\mathcal{H}_{A})$, then,
$D_{\max}(\tilde{\rho}\|\sigma)\leq\varepsilon-\log\left({1-\delta}\right),$
where $\varepsilon\geq 0,\delta\in[0,1)$ and $\tilde{\rho}:=\frac{G\rho G^{\dagger}}{\mathrm{Tr}\left[G\rho G^{\dagger}\right]}$, has the property that,
$\frac{1}{2}{\left\lVert\rho-\tilde{\rho}\right\rVert}_{1}\leq\sqrt{\delta(2-\delta)},$
where $G:=\left(e^{\varepsilon}\sigma\right)^{\frac{1}{2}}\left(e^{\varepsilon}\sigma+\left|\rho-e^{\varepsilon}\sigma\right|_{+}\right)^{-\frac{1}{2}}.$ Further, if $\operatorname{supp}(\rho)=\operatorname{supp}(\sigma),$ then, $\operatorname{supp}(\tilde{\rho})=\operatorname{supp}(\rho)=\operatorname{supp}(\sigma).$
III Blackwell’s Dominance of Quantum Mechanisms
Blackwell [9] in classical statistics provided a framework to compare two statistical experiments in terms of their informativeness. An experiment is considered more informatively dominating if it allows for better decision-making based on the observed data. This idea of dominance can be viewed from the point of view of hypothesis testing. Consider $\mathcal{E}_{1}:=\left\{P_{1},P_{2}\right\}$ and $\mathcal{E}_{2}:=\left\{Q_{1},Q_{2}\right\}$ be two classical statistical experiments, where $P_{1},P_{2}$ and $Q_{1},Q_{2}$ are probability distributions over two finite sets $\mathcal{X}$ and $\mathcal{Y}$. Consider the two following Hypothesis tests:
$H(\mathcal{E}_{1})$ $:=\begin{cases}H^{\mathcal{E}_{1}}_{0}:\mbox{A random variable }X\sim P_{1},\\ H^{\mathcal{E}_{1}}_{1}:\mbox{A random variable }X\sim P_{2}.\end{cases}$
$H(\mathcal{E}_{2})$ $:=\begin{cases}H^{\mathcal{E}_{2}}_{0}:\mbox{A random variable }Y\sim Q_{1},\\ H^{\mathcal{E}_{2}}_{1}:\mbox{A random variable }Y\sim Q_{2}.\end{cases}$
We now say that $\mathcal{E}_{2}$ is dominated by $\mathcal{E}_{1}$ (written as $\mathcal{E}_{2}\preceq_{B}\mathcal{E}_{1}$) if for any test (decision rule) $\phi_{\mathcal{E}_{2}}:\mathcal{Y}\to[0,1]$ for $H(\mathcal{E}_{2})$, there exists a test $\phi_{\mathcal{E}_{1}}:\mathcal{X}\to[0,1]$ for $H(\mathcal{E}_{1})$ such that $\phi_{\mathcal{E}_{1}}$ performs at least as well as $\phi_{\mathcal{E}_{2}}$. We formalize this notion in the following proposition.
Proposition 1 (Blackwell’s Theorem of Informative Dominance[10, Theorems $12.4.1$ & $12.4.2$]).
  1. 1.
    $\left\{Q_{1},Q_{2}\right\}\preceq_{B}\left\{P_{1},P_{2}\right\}$.
  2. 2.
    $\beta(\alpha,Q_{1},Q_{2})\geq\beta(\alpha,P_{1},P_{2})$ for all $\alpha\in[0,1]$, where $\beta(\alpha,P_{1},P_{2})$ and $\beta(\alpha,Q_{1},Q_{2})$ is the minimum type-II error of order $\alpha$ (see Definition 4) of the hypothesis tests $H(\left\{P_{1},P_{2}\right\})$ and $H(\left\{Q_{1},Q_{2}\right\})$ respectively.
  3. 3.
    $D^{\alpha}_{H}{(Q_{1}\|Q_{2})}\leq D^{\alpha}_{H}{(P_{1}\|P_{2})}$ for all $\alpha\in[0,1]$, where $D^{\alpha}_{H}{(P_{1}\|P_{2})}$ and $D^{\alpha}_{H}{(Q_{1}\|Q_{2})}$ are the classical hypothesis testing divergence of order $\alpha$ (see Definition 5) between the pair of distributions $\left\{P_{1},P_{2}\right\}$ and $\left\{Q_{1},Q_{2}\right\}$ respectively.
  4. 4.
Observe that Proposition 1 above provides a partial order $\preceq_{B}$ on experiments, i.e., we say an experiment $\mathcal{E}_{2}$ is dominated by (or less informative than) another experiment $\mathcal{E}_{1}$ if $\mathcal{E}_{2}\preceq_{B}\mathcal{E}_{1}$.
Similarly, in binary asymmetric quantum hypothesis testing, given any pair of quantum states $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, we perform a hypothesis test between $\rho$, then, the null hypothesis, and $\sigma$, the alternative hypothesis. The test is biased towards the null hypothesis, i.e., if the given state is originally $\rho$, then the test should reject the null hypothesis with a small probability, while accepting the null hypothesis with a small probability if the given state is $\sigma$. A quantum hypothesis test is typically performed using a binary POVM $\{\Lambda,\mathbb{I}-\Lambda\}$, where $\Lambda$ is the rejection operator and $\mathbb{I}-\Lambda$ is the acceptance operator. The hypothesis test can incur two types of errors, namely, the type-I error (false negative) and the type-II error (false positive). The type I error is the probability of rejecting the null hypothesis when it is true, while the type II error is the probability of accepting the null hypothesis when it is false. Under the binary POVM $\{\Lambda,\mathbb{I}-\Lambda\}$, the type-I error is given by $\mathrm{Tr}[\Lambda\rho]$ and the type-II error is given by $\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma]$. To characterize the trade-off between the type-I and type-II errors, in asymmetric quantum hypothesis testing, we fix the type-I error to a value $\alpha\in[0,1]$ and maximize the negative logarithm of type-II error over all possible POVMs $\{\Lambda,\mathbb{I}-\Lambda\}$, which gives us the following definition of quantum hypothesis testing divergence.
Definition 14 (Quantum hypothesis testing divergence [28]).
For any pair of quantum states $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, we denote its fixed point as the level of type-I error $\alpha\in[0,1]$ for which the type-II error is equal to the type-I error, i.e., $e^{-D^{\alpha}_{H}{(\rho\|\sigma)}}=\alpha$. This is the point where both hypotheses are equally likely. Similar to classical Blackwell’s ordering, we define quantum Blackwell’s order as follows,
Definition 15 (Quantum Blackwell’s order).
Given two pairs of quantum states $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$ over two finite-dimensional Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, respectively, we say that $\left\{\sigma_{1},\sigma_{2}\right\}$ is dominated by (or less informative than) $\left\{\rho_{1},\rho_{2}\right\}$ (denoted as $\left\{\sigma_{1},\sigma_{2}\right\}\preceq_{B_{Q}}\left\{\rho_{1},\rho_{2}\right\}$) if for any $\alpha\in[0,1]$, the following holds,
Lemma 1 (Quantum Blackwell’s Theorem).
Given two pairs of quantum states $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$ over two finite-dimensional Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, respectively, the following statements are equivalent,
  1. 1.
    $\left\{\sigma_{1},\sigma_{2}\right\}\preceq_{B_{Q}}\left\{\rho_{1},\rho_{2}\right\}$.
  2. 2.
    There exists a completely positive map $\mathcal{T}:\mathcal{D}(\mathcal{H}_{A})\to\mathcal{D}(\mathcal{H}_{B})$, such that $\sigma_{1}=\mathcal{T}(\rho_{1})$ and $\sigma_{2}=\mathcal{T}(\rho_{2})$.
Proof.
Before proceeding with the proof of Lemma 1, we first prove the following lemma, which will be useful in proving Lemma 1.
Lemma 2.
then,
$E_{\gamma}(\rho_{1}\|\rho_{2})\geq E_{\gamma}(\sigma_{1}\|\sigma_{2}),\forall\gamma\geq 0,$
where $E_{\gamma}(\cdot\|\cdot)$ is defined in Definition 13 for any $\gamma\geq 0$.
Proof.
See Appendix -A for the proof.
It is easy to follow that $(2)\implies(1)$ as a consequence of Definition 15 and Fact 12. We now proceed to prove $(1)\implies(2)$. From Definition 15 $(1)$ implies (28). Further, from Lemma 2 and Fact 10, it follows that, if for the pairs $\rho_{1},\rho_{2}\in\mathcal{D}(\mathcal{H}_{A})$ and $\sigma_{1},\sigma_{2}\in\mathcal{D}(\mathcal{H}_{B})$, (28) holds, then there exists a completely positive map $\mathcal{T}:\mathcal{H}_{A}\to\mathcal{H}_{B}$ such that $\sigma_{i}=\mathcal{T}(\rho_{i})$, for each $i\in\{1,2\}$. This completes the proof of Lemma 1.
The concept of Blackwell dominance for pairs of quantum states, as presented in Definition 15, is related to the framework for the statistical comparison of quantum channels developed by Buscemi [11]. In that work, an informational ordering is defined between two quantum channels, $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, based on the guessing probability of any ensemble of states passed through them. Specifically, channel $\mathcal{N}_{1}$ is considered more informative than $\mathcal{N}_{2}$ if, for every ensemble of input states, the guessing probability after passing through $\mathcal{N}_{1}$ is at least as high as that after passing through $\mathcal{N}_{2}$. This ordering is shown to be equivalent to the existence of a completely positive map that transforms the output of $\mathcal{N}_{1}$ into that of $\mathcal{N}_{2}$, mirroring the condition in Lemma 1. However, it remains an open question whether this ordering can be fully characterized by the quantum hypothesis testing divergence, as in Definition 15. Addressing this question would deepen our understanding of the relationship between quantum channel comparison and hypothesis testing.
@@ III Blackwell’s Dominance of Quantum Mechanisms
IV Characteristic Region of $(\varepsilon,\delta)$-Differentially Private Quantum Mechanisms
@@
In this section, we introduce the concept of the characteristic region of a quantum mechanism that satisfies $(\varepsilon,\delta)$-quantum differential privacy (QDP) as defined in [4, Definition $2$]. The characteristic region provides a geometric representation of the trade-offs between the privacy parameters $\varepsilon$ and $\delta$ for a given private quantum mechanism. To illustrate this concept, we consider the following definition of a pair of quantum states to be $(\varepsilon,\delta)$-DP.
Definition 16.
A pair $\rho,\sigma\in\mathcal{D}(\mathcal{H})$ is defined to be $(\varepsilon,\delta)$-DP (differentially private) for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, if every POVM measurement $0\preceq\Lambda\preceq\mathbb{I}$, the following holds,
Consider a hypothesis test between two quantum states $\rho$ and $\sigma$ over a finite dimensional Hilbert space $H$ as follows,
$H$ $:=\begin{cases}H_{0}:\mbox{Given quantum state is}~\rho,\\ H_{1}:\mbox{Given quantum state is}~\sigma.\end{cases}$
@@
Note that from Definition 15 it follows that for any two pairs of quantum states $(\rho_{1},\sigma_{1})$ and $(\rho_{2},\sigma_{2})$, we have $\{\sigma_{1},\sigma_{2}\}\preceq_{B_{Q}}\{\rho_{1},\rho_{2}\}$ if and only if $\mathcal{R}(\sigma_{1},\sigma_{2})\subseteq\mathcal{R}(\rho_{1},\rho_{2})$. Further, if $(\rho,\sigma)$ is $(\varepsilon,\delta)$-QDP, then for any rejection POVM $\Lambda$, from (29) the following constraints hold,
$\beta_{\Lambda}$ $\geq e^{-\varepsilon}(1-\delta-\alpha_{S}),$ (32)
$\beta_{\Lambda}$ $\geq 1-\delta-e^{\varepsilon}\alpha_{S},$ (33)
$\beta_{\Lambda}$ $\leq 1-e^{-\varepsilon}(\alpha_{S}-\delta),$ (34)
$\beta_{\Lambda}$ $\leq e^{\varepsilon}(1-\alpha_{S})+\delta.$ (35)
@@
1n Figure 1, we illustrate a graphical representation of the characteristic region $\mathcal{R}(\varepsilon,\delta)$ of $(\varepsilon,\delta)$-QDP.
Therefore, under the constraints mentioned in eqs.˜32, 33, 34 and 35, we define the characteristic (or operating) region of quantum $(\varepsilon,\delta)$-QDP as follows,
@@
From eqs.˜32, 33, 34 and 35, it is trivial to see that for any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ (for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$), we have $\mathcal{R}(\rho,\sigma)\subseteq\mathcal{R}(\varepsilon,\delta)$. Moreover, the characteristic region $\mathcal{R}(\varepsilon,\delta)$ is a convex set. This is because for any two positive operators $0\preceq\Lambda_{1},\Lambda_{2}\preceq\mathbb{I}$, for any $\theta\in[0,1]$, the operator $\Lambda=\theta\Lambda_{1}+(1-\theta)\Lambda_{2}$ is also a positive operator satisfying $0\preceq\Lambda\preceq\mathbb{I}$. Therefore, it can be verified that $\alpha_{\Lambda}=\theta\alpha_{\Lambda_{1}}+(1-\theta)\alpha_{\Lambda_{2}}$ and $\beta_{\Lambda}=\theta\beta_{\Lambda_{1}}+(1-\theta)\beta_{\Lambda_{2}}$, where $(\alpha_{\Lambda_{1}},\beta_{\Lambda_{1}})$ and $(\alpha_{\Lambda_{2}},\beta_{\Lambda_{2}})$ belongs to $\mathcal{R}(\varepsilon,\delta)$. Thus, as $(\alpha_{\Lambda},\beta_{\Lambda})$ also belongs to $\mathcal{R}(\varepsilon,\delta)$, this implies that $\mathcal{R}(\varepsilon,\delta)$ is convex. Further, $\mathcal{R}(\varepsilon,\delta)$ is also a closed set. We will prove this by constructing a pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})\in\mathfrak{D_{(\varepsilon,\delta)}}$, which achieves all the corner points of $\mathcal{R}(\varepsilon,\delta)$, as mentioned in Figure 1 above.
@@
Consider that $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ are two quantum states over a finite-dimensional Hilbert space $\mathcal{H}$ such that for an one-dimensional projector $|v\rangle\langle v|$ we have,
@@
Now, we notice that the cardinality of the support of the state pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ must be at least $3$ since we need two more corner points to be achieved.
@@
Finally, consider the one-dimensional projector $|z\rangle\langle z|=\mathbb{I}-|v\rangle\langle v|-|x\rangle\langle x|-|y\rangle\langle y|$ which is perpendicular to $|v\rangle\langle v|$, $|x\rangle\langle x|$ and $|y\rangle\langle y|$. Then, from eqs.˜37, 38, 39, 40, 45 and 46 we have,
@@
Remark 1.
  • It is important to note that in (43) and (44), if we had chosen the RHS to be $(\frac{1-\delta}{1+e^{\varepsilon}}$ and $\frac{1-\delta}{1+e^{\varepsilon}}$ we would get another pair of quantum states $(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ which also achieves the extremal points of the characteristic region $\mathcal{R}(\varepsilon,\delta)$ for any given $\varepsilon\geq 0$ and $\delta\in[0,1]$. This pair is given as follows,
In Figure 2 below, we illustrate a graphical representation of the characteristic region $\mathcal{R}(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ along with the extremal points achieved by the pair $(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$.
@@
V Differential Privacy : A Quantum Hypothesis Testing Perspective
V-A Main characterization
In this section, we consider differential privacy from a quantum hypothesis testing perspective. The main goal of quantum differential privacy is to ensure that a quantum mechanism produces output states that are nearly indistinguishable when applied to neighboring quantum inputs. Formally, for two neighboring quantum inputs producing output states $\rho,\sigma$, the distinguishability between $\rho$ and $\sigma$ must be limited by the privacy parameters. Intuitively, this means that even if an investigator (adversary) knows the mechanism and observes its output, the investigator should not be able to reliably infer which quantum input (respondent) generated it. More formally, with high probability, no hypothesis test—regardless of the investigator’s strategy—can reliably infer the respondent’s individual contribution from the output.
Further, it is important to observe that given two pairs of quantum states $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$, if for any $\alpha\in[0,1]$ the hypothesis testing divergence satisfies if $D^{\alpha}_{H}{(\rho_{1}\|\rho_{2})}\geq D^{\alpha}_{H}{(\sigma_{1}\|\sigma_{2})}$, then $(\sigma_{1},\sigma_{2})$ is harder to distinguish than $(\rho_{1},\rho_{2})$ for all possible values of the Type-I error parameter $\alpha$. Intuitively, this indicates that the states $\rho_{1}$ and $\rho_{2}$ are more distinguishable—or equivalently, more well-separated—than the states $\sigma_{1}$ and $\sigma_{2}$. This observation naturally leads to the following result.

### v3

@@ II Notations, Definitions and Facts
III Characteristic Region of $(\varepsilon,\delta)$-Differentially Private Quantum Mechanisms
@@
In this section, we introduce the concept of the characteristic region of a quantum mechanism that satisfies $(\varepsilon,\delta)$-quantum differential privacy (QDP) as defined in [36, Definition $2$]. The characteristic region provides a geometric representation of the trade-offs between the privacy parameters $\varepsilon$ and $\delta$ for a given private quantum mechanism. To illustrate this concept, consider a hypothesis test between two quantum states $\rho$ and $\sigma$ over a finite dimensional Hilbert space $H$ as follows,
$H$ $:=\begin{cases}H_{0}:\mbox{Given quantum state is}~\rho,\\ H_{1}:\mbox{Given quantum state is}~\sigma.\end{cases}$ (26)
@@
Further, if $(\rho,\sigma)$ is $(\varepsilon,\delta)$-QDP, then for any rejection POVM $\Lambda$, from (1) the following constraints hold,
$\beta_{\Lambda}$ $\geq e^{-\varepsilon}(1-\delta-\alpha_{\Lambda}),$ (28)
$\beta_{\Lambda}$ $\geq 1-\delta-e^{\varepsilon}\alpha_{\Lambda},$ (29)
$\beta_{\Lambda}$ $\leq 1-e^{-\varepsilon}(\alpha_{\Lambda}-\delta),$ (30)
$\beta_{\Lambda}$ $\leq e^{\varepsilon}(1-\alpha_{\Lambda})+\delta.$ (31)
@@
In Figure 1, we illustrate a graphical representation of the characteristic region $\mathcal{R}(\varepsilon,\delta)$ of $(\varepsilon,\delta)$-QDP.
Therefore, under the constraints mentioned in eqs.˜28, 29, 30 and 31, we define the characteristic (or operating) region of quantum $(\varepsilon,\delta)$-QDP as follows,
@@
From eqs.˜28, 29, 30 and 31, it is trivial to see that for any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ (for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$), we have $\mathcal{R}(\rho,\sigma)\subseteq\mathcal{R}(\varepsilon,\delta)$. Moreover, the characteristic region $\mathcal{R}(\varepsilon,\delta)$ is a convex set. This is because for any two positive operators $0\preceq\Lambda_{1},\Lambda_{2}\preceq\mathbb{I}$, for any $\theta\in[0,1]$, the operator $\Lambda=\theta\Lambda_{1}+(1-\theta)\Lambda_{2}$ is also a positive operator satisfying $0\preceq\Lambda\preceq\mathbb{I}$. Therefore, it can be verified that $\alpha_{\Lambda}=\theta\alpha_{\Lambda_{1}}+(1-\theta)\alpha_{\Lambda_{2}}$ and $\beta_{\Lambda}=\theta\beta_{\Lambda_{1}}+(1-\theta)\beta_{\Lambda_{2}}$, where $(\alpha_{\Lambda_{1}},\beta_{\Lambda_{1}})$ and $(\alpha_{\Lambda_{2}},\beta_{\Lambda_{2}})$ belongs to $\mathcal{R}(\varepsilon,\delta)$. Thus, as $(\alpha_{\Lambda},\beta_{\Lambda})$ also belongs to $\mathcal{R}(\varepsilon,\delta)$, this implies that $\mathcal{R}(\varepsilon,\delta)$ is convex. Further, $\mathcal{R}(\varepsilon,\delta)$ is also a closed set. We will prove this by constructing a pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})\in\mathfrak{D_{(\varepsilon,\delta)}}$, which achieves all the corner points of $\mathcal{R}(\varepsilon,\delta)$, as mentioned in Figure 1 above.
@@
Consider a pair $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ over a finite-dimensional Hilbert space $\mathcal{H}$ such that for a one-dimensional projector $|v\rangle\langle v|$,
@@
Now, we notice that the cardinality of the support of the state pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ must be at least $3$, since we need two more corner points to be achieved.
@@
Finally, consider the one-dimensional projector $|z\rangle\langle z|=\mathbb{I}-|v\rangle\langle v|-|x\rangle\langle x|-|y\rangle\langle y|$ which is perpendicular to $|v\rangle\langle v|$, $|x\rangle\langle x|$ and $|y\rangle\langle y|$. Then, from eqs.˜33, 34, 35, 36, 41 and 42 we have,
@@
Remark 1.
  • It is important to note that in (39) and (40), if we had chosen the RHS to be $\frac{1-\delta}{1+e^{\varepsilon}}$ and $\frac{1-\delta}{1+e^{\varepsilon}}$ respectively, we would get another pair of quantum states $(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ which also achieves the extremal points of the characteristic region $\mathcal{R}(\varepsilon,\delta)$ for any given $\varepsilon\geq 0$ and $\delta\in[0,1]$. This pair is given as follows,
In Figure 2 above, we illustrate a graphical representation of the characteristic region $\mathcal{R}(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ along with the extremal points achieved by the pair $(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$.
Remark 2.
In [5], the authors obtained a figure similar to Figure 1 by plotting type-1 and type-2 errors for a pair of states $\{\rho_{1},\rho_{2}\}$. In contrast, Figure 1 is with respect to all pairs $\{\rho_{1},\rho_{2}\}$ which are $(\varepsilon,\delta)$-DP and therefore it helps to characterize the worst-case (most informative) $(\varepsilon,\delta)$-DP pair of quantum states.
IV Quantum Information Ordering and Differential Privacy
IV-A Quantum ordering of informativeness
In [4, 3], Blackwell introduced a notion of an ordering of classical informativeness between two pairs of distributions $\{P_{1},P_{2}\}$ and $\{Q_{1},Q_{2}\}$. In particular, he proved the following proposition,
Proposition 1 (Blackwell’s Theorem of Informativeness Ordering [3, Theorems $12.2.2$ & $12.4.2$]).
  1. 1.
    $\left\{P_{1},P_{2}\right\}\succeq_{B}\left\{Q_{1},Q_{2}\right\}$.
  2. 2.
    $D^{\alpha}_{H}{(P_{1}\|P_{2})}\geq D^{\alpha}_{H}{(Q_{1}\|Q_{2})}$ for all $\alpha\in[0,1]$, where $D^{\alpha}_{H}{(P_{1}\|P_{2})}$ and $D^{\alpha}_{H}{(Q_{1}\|Q_{2})}$ are the classical hypothesis testing divergence of order $\alpha$ (see Definition 3) between the pair of distributions $\left\{P_{1},P_{2}\right\}$ and $\left\{Q_{1},Q_{2}\right\}$ respectively.
  3. 3.
Analogous to Blackwell’s order of classical informativeness (as mentioned in Proposition 1), we discuss a quantum version of it, which provides an ordering of quantum statistical experiments based on their distinguishability. Using the quantum hypothesis testing divergence as the underlying measure, we formalize when one pair of quantum states can be considered more informative than another. We further establish that this dominance relation implies a corresponding ordering across all quantum $f$-divergences, thereby linking hypothesis testing quantities to a broad class of information-theoretic measures.
We now make the following definition.
Definition 17 (Quantum information order).
Consider two pairs of quantum states $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$. We say that the pair $\left(\rho_{1},\rho_{2}\right)$ dominates (or, is more informative than) the pair $\left(\sigma_{1},\sigma_{2}\right)$ (denoted as $\left\{\rho_{1},\rho_{2}\right\}\succeq_{Q}\left\{\sigma_{1},\sigma_{2}\right\}$), if for all $\alpha\in[0,1]$, the following holds,
Intuitively, the order $\left\{\rho_{1},\rho_{2}\right\}\succeq_{Q}\left\{\sigma_{1},\sigma_{2}\right\}$ indicates that the states $\rho_{1}$ and $\rho_{2}$ are more distinguishable as compared to the states $\sigma_{1}$ and $\sigma_{2}$.
[…]
However, we can circumvent this issue by directly showing that if for any two pairs of quantum states $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$, (51) holds, then, every quantum $f$-divergence (see Definition 10) between the former pair dominates that of the latter pair. We formalize this claim in the Theorem below.